an elephant of mass 1500kg stands on its four legs each of area 0.012m^2, calculate the pressure exerted by each leg of the elephant on the ground.(g=10m/s)?


1
Expert's answer
2022-02-20T08:20:06-0500

Given:

"m=1500\:\rm kg"

"A=4*0.012\:\rm m^2=0.048\: m^2"


The pressure

"P=mg/A\\
=\frac{1500\:\rm kg*10\: N/kg}{0.048\:\rm m^2}=312500\:\rm Pa"